# D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts

### v5

@@ VII Defense-Driven Discrimination of Distributed Concepts (D4C)
VII-A Negative examples are not merely intended to confound the discriminator
@@
Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern where positive examples yield high probabilities when substituted into the case’s problem statement, while negative examples yield low probabilities. We can view this pattern in the following way: the discriminator constructs a solution distribution for the cases presented in abstract reasoning problems. Within this distribution, positive examples are assigned higher probabilities, whereas negative examples are assigned lower ones. The endeavor to train the discriminator can be conceptualized as a process of optimizing and fine-tuning the solution distribution, with the ultimate objective of aligning it more closely with the underlying distribution of correct solutions for the given cases. Using positive and negative examples to estimate a Gaussian distribution through the cross-entropy loss function can be seen as leveraging them to optimize the following function:
@@
Where $x$ represents the positive or negative samples, $\mu$ and $\sigma$ represent the mean and standard deviation of the Gaussian distribution separately. It can be observed that when positive instances are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the mean squared loss between sample $x$ and the mean of the correct solution $\mu$. Conversely, when negative instances are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$ of the respective segment.
@@
VII-B High-quality negative examples lead to more rigorous concept boundaries
@@
The D2C methodology proposed in this study has established a novel problem boundary for deep networks by means of resampling negative instances and incorporating a fresh loss function. Experimental findings tentatively indicate that this boundary is better suited for scenarios characterized by the absence of conceptual supervision signals, as compared to the boundaries delineated by the original negative instances in Bongard-Logo. Nonetheless, the precise tightness of the new boundary remains indeterminate.
@@
In essence, D2C introduces a new variance to the solution distribution of the network, the magnitude of which relative to the variance of the correct solution distribution is indeterminate. In the context of the Bongard-Logo challenge, the variance provided by D2C appears closer to that of the correct solution distribution in the absence of conceptual supervision signals, as compared to the variance described by the original negative instances.
@@
Evidently, the efficacy of D2C is contingent upon specific dataset attributes. As anticipated, empirical evidence suggests that this approach may be insufficient for delineating precise boundaries within the framework of the RPM problem. Contrary to the Bongard-Logo paradigm, RPM problems postulate a shared set of core concepts across distinct instances, ensuring discreteness and stability. This shared nature precludes the utility of methods like D2C in addressing RPM challenges. Specifically, image properties within RPM instances are rigidly governed by definitions of “rules” or “progressive patterns,” maintaining consistent attributions across all instances.
@@
The Bongard-Logo problem exhibits varying attributions of image properties due to differing clustering bases among instances, thereby augmenting problem complexity and difficulty. Incorporating positive instances from other cases as negative examples can, to some extent, delineate conceptual boundaries between distinct cases, adhering to a cautious approach aimed at minimizing conceptual infringement. However, this strategy proves inapplicable in addressing challenges presented by RAVEN and PGM due to substantial disparities across multiple dimensions.
@@
In summary, the success of D2C in addressing the Bongard-Logo problem can be attributed to its indirect enhancement of the quality of negative instances within Bongard-Logo cases. Conversely, its limitations in tackling the RPM problem can also be ascribed to its insignificant improvement in the quality of negative examples within RPM scenarios. To address the limitations inherent in the D2C approach, this paper introduces the D4C (Defense-Driven Discrimination of Distributed Concepts) method. The underlying idea of D4C is to generate higher-quality negative instances and samples within abstract reasoning problems. This approach aims to delineate a more appropriate problem boundary for the discriminator as compared to the original problem setup in abstract reasoning. By leveraging novel mechanisms and strategies, D4C aims to deliver more precise and reliable definitions of conceptual boundaries and foster further advancements in the realm of abstract reasoning.
@@ VII-C Impelmention of D4C method
Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC and Triple-CFN, achieving notable successes. These discriminators seamlessly integrate with the D2C and D3C methodologies proposed in this study, and pertinent experiments have been conducted to explore their combinations. These discriminators excel in accurately identifying and classifying Bongard-Logo instances, significantly contributing to the advancement of research in this domain.
@@
For the generator in Triple-CFN, it is essential to attenuate the generated representations. Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group. After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to attenuate the confusing representation vector. This attenuation process is necessary due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation. This makes it easier for a generator based on the Transformer encoder to generate realistic confusing representations, potentially causing Triple-CFN to collapse. Therefore, we have designed the aforementioned attenuation process. The alignment between the generator and Triple-CFN is illustrated in Figure 17. In this figure, “$f_{\theta}(z_{ij}|x_{ij})$” represents the encoder in Triple-CFN.
@@
However, Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix. To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations. Any inconsistencies or irregularities identified within these representations will be flagged as negative instances, indicating incorrect responses. Through this rigorous process, Lico-Net strives to accurately identify and reject perplexing representations, thereby augmenting its overall precision and reliability.
@@ VIII-A Experiment on RAVEN
Initially, we conducted experiments with Lico-Net on both RAVEN[15] and I-RAVEN[37] datasets to demonstrate its inherent superiority. To ensure a fair comparison between excellent models, all experimental settings and equipment in this study were kept consistent with those of RS-Tran and Triple-CFN, including data volume, optimizer parameters, data augmentation methods (such as reducing the RAVEN image size to $80\times 80$), batch size, and other relevant hyperparameters.
@@
Triple-CFN[32] 99.6/99.8 100.0/100.0 99.7/99.8 98.8/99.4 99.9/100.0 99.9/100.0 99.9/99.9 99.2/99.2
Lico-Net 99.7/99.8 100.0/100.0 99.7/99.8 98.9/99.5 99.9/100.0 99.9/100.0 99.9/99.9 99.4/99.6
@@ VIII-B Experiment on Bongard-Logo
Consequently, we conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN as backbones for D2C, D3C, and D4C and using PMoC as backbone for D2C and D4C. The experimental results are presented in the table III. In the experiment with D3C, we conducted an additional experiment on its lightweight version, D3C-cos, to verify that replacing D3C with D3C-cos would not significantly affect performance. As can be seen in the table III, we did not conduct an experiment with Triple-CFN+D3C due to the large model size of Triple-CFN. This made it impractical to deploy the model using conceptual distribution and logical optimization with the Sinkhorn distance on our existing experimental equipment. This was the driving force behind proposing D3C-cos as an alternative to D3C in this paper. In addition, the combination of D4C and PMoC+Straw Pose-Transformer has not been implemented due to limitations in computing power.
@@
PMoC 92.0 92.6 97.7 78.3 75.0
PMoC+Straw Pose-Transformer 96.2 96.0 98.5 78.5 75.3
Triple-CFN 93.2 92.0 99.2 80.8 79.1
ResNet50 92.0 88.9 98.1 76.1 76.0
ResNet50+D2C 92.4 90.7 99.0 77.0 77.5
ResNet50+D3C 91.2 91.0 98.5 75.5 76.9
ResNet50+D2C+D3C 96.6 93.2 99.6 78.8 78.8
ResNet50+D3C-cos 95.2 90.2 98.8 79.0 78.1
ResNet50+D2C+D3C-cos 96.0 92.3 98.8 79.8 79.1
Triple-CFN+D2C 96.5 94.8 99.4 79.5 81.6
Triple-CFN+D3C-cos 97.0 96.2 99.4 82.0 79.7
Triple-CFN+D2C+D3C-cos 97.2 96.2 99.4 82.8 82.7
Triple-CFN+D4C 97.7 93.5 99.6 82.8 83.1
PMoC+D2C 93.9 95.0 98.1 78.5 78.4
PMoC+Straw Pose-Transformer+D2C 98.7 96.8 98.9 81.3 75.6
@@
Upon observing the results presented in the table III, we found that D2C, D3C, and D3C-cos made considerable contributions in the Bongard-Logo task, with D3C-cos emerging as an effective alternative to D3C. Additionally, D4C achieved remarkable results, despite not obtaining the best performance across all test data, possibly due to the inherent limitations of the PMoC model. However, compared to its backbone, PMoC, D4C demonstrated significant improvement and contribution, validating the effectiveness of defining tighter problem boundaries for the model. This approach warrants further investigation and development in future research.
@@ VIII-C Experiment on PGM
In this paper, we conducted experiments related to Lico-Net on PGM. And the results are recorded in Tabel IV.
TABLE IV: Reasoning Accuracies of models on PGM.
@@
D4C (with Lico-Net as the backbone) 98.6(98.64)
D4C glove-train(with Lico-Net as the backbone) 98.7(98.73)
@@
Triple-CFN+Re-space layer+D3C 99.4(99.36)
Triple-CFN+Re-space layer+D3C-cos 99.4(99.37)
Meta Lico-Net 99.3(99.30)
@@
In the experimental results recorded in Tabel IV, we can observe various phenomena. Lico-Net, in essence, addresses the imperfections within the Triple-CFN framework. In Triple-CFN, concepts and minimal reasoning units serve as queries and key-value pairs, respectively, for computing cross-attention outcomes. However, a fundamental flaw in its design lies in the logical hierarchy mismatch between concepts and minimal reasoning units. Specifically, concepts often require the combined expression of multiple minimal reasoning units. Despite this shortcoming, Triple-CFN remains a remarkable model. Lico-Net, born from the rectification of these imperfections, has achieved a little improvements in accuracy.
The combination of Lico-Net and D2C, and the combination of Triple-CFN and D2C, did not bring significant improvement, with the enhancement only reflected in the percentile of reasoning accuracy. This phenomenon was anticipated in this paper, as we believe it is due to the difference between RPM and Bongard-Logo problems. In RPM problems, different cases share the same concept space and attribute recognition method. However, the combination of D4C and D4C glove-train significantly improved Lico-Net’s performance. Additionally, the glove-shaped generator proved to be more suitable for generating negative representations of RPM problems, challenging Lico-Net and enabling it to achieve further refinement.
The combination of Lico-Net and D3C is noteworthy. To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted experiments using Meta Lico-Net. Unlike “Lico-Net+D3C”, which encodes the description of image progressive patterns into distributions using a Transformer-Encoder, Meta Lico-Net encodes them into vectors. To simplify, instead of dividing the encoded representation into mu and logvar, we directly utilize the representation as a vector. Furthermore, the cosine distance between progressive pattern vectors $\{\overline{P}_{nm}\}$ and the vectorized pattern description in Lico-Net is constrained using infoNCE. In fact, the process mentioned above is an emulation of how Triple-CFN evolved into Meta Triple-CFN. The structure of Meta Lico-Net is illustrate in figure 19. The intention behind establishing Meta Lico-Net and Triple-CFN+D3C is to demonstrate that, compared to vectors, distributions are more suitable for describing abstract concepts. To further validate the theme of D3C, we applied both D3C and D3C-cos to Triple-CFN and Lico-Net. Through experimentation, we found that D3C and D3C-cos brought improvements to Triple-CFN, elevating its performance to the level of Lico-Net.
Figure 19: Feedforward process of Meta Lico-Net
TABLE V: Progressive Pattern Reasoning Accuracies and of models on PGM.
@@
The D3C’s reasoning accuracys of progressive pattern are recorded in Table V, highlighting the strong interpretability prior to training possessed by both D3C and Lico-Net. This exceeds that of Rs-Tran and its derivative, Tran-clip. It must be acknowledged that the experiment of applying D3C, which is based on the Sinkhorn distance, to Lico-Net was not conducted to its fullest extent in this paper. Once the experimental results of Lico-Net+D3C were sufficient to illustrate our point, we halted the model training. This was due to the significant computational resources required, even with the experimental equipment used in this paper: four A100s graphics cards. The training involved numerous epochs and tedious individual training epochs. This was the intended purpose behind the design of D3C-cos as a replacement for D3C, as the power of Zeus needs to be conveyed by Hermes.
VIII-D Ablation experiment on D2C
In the final of this section, we devised an experiment to validate the D2C approach. Utilizing a 500-step DDPM model, we generated a substantial quantity of Bongard-Logo data, specifically 3,180 Bongard-Logo images. According to the DDPM theory, these images belong to the same distribution as the real data, albeit with unknown semantics and concepts. During the training of ResNet50 and Triple-CFN, for each feedforward pass, we introduced 14 randomly selected DDPM-generated images as auxiliary group examples for the Bongard-Logo cases. The experimental results were recorded in “D2C-” item of the Table VI.
TABLE VI: Reasoning Accuracies of models on Bongard-logo.

### v9

@@ VII Defense-Driven Discrimination of Distributed Concepts (D4C)
VII-A Negative samples are not merely intended to confound the discriminator
@@
Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern where positive examples yield high probabilities when substituted into the case’s problem statement, while negative examples yield low probabilities. In a nutshell, they are models with probabilistic labeling scheme. We can view this pattern in the following way: the discriminator constructs a solution distribution for the cases presented in abstract reasoning problems. Within this distribution, positive examples are assigned higher probabilities, whereas negative examples are assigned lower ones. The endeavor to train the discriminator can be conceptualized as a process of optimizing and fine-tuning the solution distribution, with the ultimate objective of aligning it more closely with the underlying distribution of correct solutions for the given cases. Using positive and negative examples to estimate a Gaussian distribution through the cross-entropy loss function can be seen as leveraging them to optimize the following function:
@@
Where $x$ represents the positive or negative samples, $\mu$ and $\sigma$ represent the mean and standard deviation of the Gaussian distribution separately. It can be observed that when positive samples are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the mean squared loss between sample $x$ and the mean of the correct solution $\mu$. Conversely, when negative samples are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$ of the respective segment.
@@
VII-B High-quality negative samples lead to more rigorous concept boundaries
@@
The D2C methodology proposed in this study has established a novel problem boundary for deep networks by means of resampling negative samples and incorporating a fresh loss function. Experimental findings tentatively indicate that this boundary is better suited for scenarios characterized by the absence of concept supervision signals, as compared to the boundaries delineated by the original negative samples in Bongard-Logo. Nonetheless, the precise tightness of the new boundary remains indeterminate.
@@
In essence, D2C introduces a new variance to the solution distribution of the network, the magnitude of which relative to the variance of the correct solution distribution is indeterminate. In the context of the Bongard-Logo challenge, the variance provided by D2C appears closer to that of the correct solution distribution in the absence of concept supervision signals, as compared to the variance described by the original negative samples.
@@
Evidently, the efficacy of D2C is contingent upon specific dataset attributes. As anticipated, empirical evidence suggests that this approach may be insufficient for delineating precise boundaries within the framework of the RPM problem. Contrary to the Bongard-Logo paradigm, RPM problems postulate a shared set of core concepts across distinct cases, ensuring discreteness and stability. This shared nature precludes the utility of methods like D2C in addressing RPM challenges. Specifically, image properties within RPM cases are rigidly governed by definitions of “rules” or “progressive patterns,” maintaining consistent attributions across all cases.
@@
The Bongard-Logo problem exhibits varying identifications and determinations of image attribute due to differing clustering bases among cases, thereby augmenting problem complexity and difficulty. Incorporating positive samples from other cases as negative examples can, to some extent, delineate concept boundaries between distinct cases, adhering to a cautious approach aimed at minimizing concept infringement. However, this strategy proves inapplicable in addressing challenges presented by RAVEN and PGM due to substantial disparities across multiple dimensions.
@@
In summary, the success of D2C in addressing the Bongard-Logo problem can be attributed to its indirect enhancement of the quality of negative samples within Bongard-Logo cases. Conversely, its limitations in tackling the RPM problem can also be ascribed to its insignificant improvement in the quality of negative examples within RPM scenarios. To address the limitations inherent in the D2C approach, this paper introduces the D4C (Defense-Driven Discrimination of Distributed Concepts) method. The underlying idea of D4C is to generate higher-quality negative samples and samples within abstract reasoning problems. This approach aims to delineate a more appropriate problem boundary for the discriminator as compared to the original problem setup in abstract reasoning. By leveraging novel mechanisms and strategies, D4C aims to deliver more precise and reliable definitions of concept boundaries and foster further advancements in the realm of abstract reasoning.
@@ VII-C Impelmention of D4C method
Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC and Triple-CFN, achieving notable successes. These discriminators seamlessly integrate with the D2C and D3C methodologies proposed in this study, and pertinent experiments have been conducted to explore their combinations. These discriminators excel in accurately identifying and classifying Bongard-Logo cases, significantly contributing to the advancement of research in this domain.
@@
For the generator in Triple-CFN, it is essential to decay the generated representations. Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group. After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to decay the confusing representation. This decay process is necessary due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation. This mechanism makes it easier for a generator based on the Transformer encoder to generate realistic confusing representations, potentially causing Triple-CFN to collapse. Therefore, we have designed the aforementioned decay process. The alignment between the generator and Triple-CFN is illustrated in Figure 17. In this figure, “$f_{\theta}(z_{ij}|x_{ij})$” represents the encoder in Triple-CFN.
@@
However, Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix. To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations. Any inconsistencies or irregularities identified within these representations will be flagged as negative samples, indicating incorrect responses. Through this rigorous process, Lico-Net strives to accurately identify and reject perplexing representations, thereby augmenting its overall precision and reliability.
@@ VIII-A Experiment on RAVEN
Initially, we conducted experiments with Lico-Net on both RAVEN[15] and I-RAVEN[37] datasets to demonstrate its inherent potential. To ensure a fair comparison between excellent models, all experimental settings and equipment in this study were kept consistent with those of RS-Tran [28] and Triple-CFN [32], including data volume, optimizer parameters, data augmentation methods (such as reducing the RAVEN image size to $80\times 80$), batch size, and other relevant hyperparameters.
@@
Triple-CFN[32] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
Triple-CFN+Re-space[32] 99.4/99.6 100.0/100.0 99.7/99.8 98.0/99.1 99.9/100.0 99.9/100.0 99.9/99.9 98.5/99.0
Lico-Net 99.5/99.7 100.0/100.0 99.7/99.8 98.2/99.2 99.9/100.0 99.9/100.0 99.9/99.9 98.8/99.2
An interesting observation from the Table II indicates a consensus among existing excellent RPM solvers, including Lico-Net. Specifically, these solvers exhibit differing values of reasoning accuracy when tackling the RAVEN [15] and I-RAVEN [37] datasets. The I-RAVEN dataset was designed to address certain limitations in the option pool configuration of RAVEN by introducing improvements to its negative example setup. These enhancements include, but are not limited to, increasing the number of attributes that are altered in the negative examples and augmenting the magnitude of attribute deviations within them. However, from a human perspective, such additional setups do not alter the difficulty level of reasoning, and thus do not make I-RAVEN any harder or easier compared to RAVEN. This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the superior quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN. Such phenomenons echo the D4C’s spirit.
@@ VIII-B Experiment on Bongard-Logo
We conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN [32] as backbones for D2C, D3C(-cos), and D4C and using PMoC [31] as backbone for D2C and D4C. PMoC, a probabilistic labeling model, does not meet the requirements of D3C(-cos), as D3C(-cos) demands a backbone in the form of $f(z|x)$. The experimental results are presented in the table III.
Upon observing the results presented in the table III, we found that D2C, D3C, and D3C-cos made considerable contributions in the Bongard-Logo task. As can be seen in the table III, replacing D3C with D3C-cos to integrate with ResNet50 would not significantly affect performance. We argue that, from a results-oriented perspective, D3C-cos can be considered as a effective alternative to D3C. Notably, we did not conduct an experiment with Triple-CFN+D3C and Triple-CFN+D2C+D3C. This is because the large number of Triple-CFN parameters makes integrating the D3C approach impractical on our current experimental equipment. This was the driving force that led us to propose D3C-cos as an alternative to D3C in this paper.
@@
PMoC 92.0 90.6 97.7 77.3 76.0
Triple-CFN 93.2 92.0 98.2 78.0 78.1
ResNet50 (infoNCE baseline) 92.0 88.9 98.0 76.1 76.0
ResNet50+D2C 92.6 90.5 98.2 77.0 77.2
ResNet50+D3C 92.3 90.6 98.3 76.5 76.8
ResNet50+D2C+D3C 93.0 91.0 98.5 77.5 77.5
ResNet50+D3C-cos 92.1 90.2 98.2 76.9 76.4
ResNet50+D2C+D3C-cos 92.9 90.8 98.4 77.5 77.4
ResNet50+D4C 93.2 90.8 98.5 77.2 77.3
Triple-CFN+D2C 93.9 92.7 98.5 78.2 78.5
Triple-CFN+D3C-cos 93.7 92.5 98.3 78.2 78.2
Triple-CFN+D2C+D3C-cos 94.1 93.0 98.5 78.2 78.6
Triple-CFN+D4C 94.3 92.8 98.8 78.7 79.0
PMoC+D2C 92.5 91.2 98.0 77.7 76.7
PMoC+D4C 93.0 92.6 98.2 78.0 77.5
Additionally, D4C has imparted varying degrees of improvement to Triple-CFN, ResNet50, and PMoC, with its effectiveness rivaling that of the combined application of D2C and D3C. However, D4C did not yield an ideal improvement on the FF concept for the Bongard-logo problem when applied to Triple-CFN and ResNet50. This study hypothesizes that the suboptimal performance may be attributed to the fact that D4C was specifically designed and tailored for discriminative models with a probabilistic labeling scheme. Conversely, Triple-CFN and ResNet50 do not inherently adhere to such a framework. It has been mentioned that when integrating D4C with Triple-CFN and ResNet50, a decay mechanism was incorporated. This mechanism is a compromise made by D4C to accommodate the non-probabilistic labeling scheme of these models. From the perspective of experimental results, this decay mechanism does not appear to be a fully developed mechanism. Exploring more advanced designs to facilitate the adaptation of D4C to non-probabilistic labeling models, such as Triple-CFN, deserves attention as future work. In contrast to Triple-CFN, PMoC [31] serves as an intuitive probabilistic labeling model for the Bongard-Logo problem. Correspondingly, D4C provides a more significant enhancement when applied to PMoC, owing to its scheme.
D4C demonstrated improvement and contribution, validating the effectiveness of defining tighter problem boundaries for the model with a probabilistic labeling scheme. To integrate D4C with Triple-CFN, we devised a representation decay mechanism that, despite its imperfections, is sufficiently effective. This approach warrants further investigation and development in future research.
Finally, this paper notes that augmentation experiments were conducted for both PMoC [31] and Triple-CFN [32]. These experiments involved expanding the data volume of the Bongard-Logo dataset using data augmentation techniques, specifically including rotation and flipping. It should be emphasized that these data augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [31]. Rather, the aim of these experiments was to investigate the potential of the models with a larger dataset. Since PMoC posits that Transformer-based visual processing models require substantial data as a foundation, a point also echoed by ViT [17] and RS-Tran [28]. Furthermore, both PMoC and Triple-CFN agree that the utilization of these data augmentation methods serves to enrich the expressions of concepts at the pixel level, thereby challenging the network’s abilities to summarize, derive, accommodate, and invert concepts [31, 32]. This paper conducted experiments to verify the effectiveness of D4C with the expanded dataset. The results are recorded in Table IV.
TABLE IV: Reasoning Accuracies of D4C on Augmented Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
PMoC 94.5 92.6 98.0 78.3 76.5
Triple-CFN 94.9 93.0 99.2 80.8 79.1
@@
Triple-CFN+D4C 97.7 94.8 99.6 82.8 83.1
The results presented in Table IV demonstrate that the integration of D4C into PMoC, which is under a probabilistic labeling scheme, has led to a significant enhancement in performance. The increase in data volume has enabled D4C to further narrow the gap between PMoC and Triple-CFN.
@@ VIII-C Experiment on PGM
In this paper, we conducted experiments related to Lico-Net on PGM. And the results are recorded in Tabel V.
TABLE V: Reasoning Accuracies of models on PGM.
@@
D4C (with Lico-Net as the backbone) 98.7(98.70)
D4C glove-train(with Lico-Net as the backbone) 98.8(98.82)
RS-Tran+Tranclip[28] 99.0
Meta Triple-CFN[32] 98.4
@@
Lico-Net+D3C- 99.3(99.30)
@@
From the experimental results presented in Table V, several notable phenomena emerge. In Triple-CFN, both progressive patterns and minimal reasoning units are positioned at the same logical level [32]. However, a significant design flaw arises from the mismatch in the logical hierarchy between these patterns and units. More specifically, progressive patterns often require the combined expression of multiple minimal reasoning units. Despite this limitation, Triple-CFN still stands as an impressive model. Naturally, leveraging the inductive bias inherent to the RPM problem as the guideline in its design, Lico-Net has achieved even more remarkable results than Triple-CFN in the RPM challenges.
The combination of Lico-Net and D2C, and the combination of Triple-CFN and D2C, did not bring significant improvement, with the enhancement only reflected in the percentile of reasoning accuracy. This phenomenon was anticipated in this paper, as we believe it is due to the difference between RPM and Bongard-Logo problems. In RPM problems, different cases share the same concept space and attribute recognition method. The D2C approach, designed to address issues of concept interference and erroneous coupling among cases, is not applicable to RPM problems.
However, the combination of D4C and D4C glove-train significantly improved Lico-Net’s performance. Additionally, the glove-shaped generator proved to be more suitable for generating negative representations of RPM problems for Lico-Net and enabling it to achieve further refinement. The core innovation of this paper, D4C, has yielded promising results on both the RPM and Bongard-Logo problems, demonstrating the significant contribution of this work.
The combination of Lico-Net and D3C is noteworthy. To validate the theme of D3C, we applied both D3C and D3C-cos to Lico-Net. Through experimentation, we found that D3C and D3C-cos brought improvements to Lico-Net. To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted ablation experiments using “Lico-Net+D3C-”. Unlike “Lico-Net+D3C”, which encodes the description of image progressive patterns into distributions using a Transformer-Encoder, “Lico-Net+D3C-” encodes them into vectors. To simplify, instead of dividing the encoded representation into mu and logvar, we directly utilize the representation as a vector. Furthermore, the cosine distance between progressive pattern vectors $\{\overline{P}_{nm}\}$ and the vectorized pattern description in Lico-Net is constrained using infoNCE. In essence, the Lico-Net+D3C- can be interpreted as fixing the variance of the distributions established for all concepts in Lico-Net+D3C-cos to 0. The structure of Lico-Net+D3C- is illustrate in figure 19.
Figure 19: Feedforward process of “Lico-Net+D3C-
TABLE VI: Progressive Pattern Reasoning Accuracies and of models on PGM.
@@
The D3C’s reasoning accuracys of progressive pattern are recorded in Table VI, highlighting the interpretability expressed by Lico-Net+D3C(-cos). It must be acknowledged that the experiment of applying D3C, which is based on the Sinkhorn distance, to Lico-Net was not conducted to its fullest extent in this paper. Once the experimental results of Lico-Net+D3C were sufficient to illustrate our point, we halted the model training. This was due to the significant computational resources required, even with the experimental equipment used in this paper: four A100s graphics cards. The training involved numerous epochs and tedious individual training epochs. This was the intended purpose behind the design of D3C-cos as a replacement for D3C, as the power of Zeus needs to be conveyed by Hermes.
Despite achieving remarkable results in terms of reasoning accuracy and interpretability of progressive patterns with Lico-Net+D3C or D3C-cos, it is noteworthy that the improvement over previous works involving Meta data as additional auxiliary labels is not particularly striking. The possible reason for this is that the introduction of Meta data significantly reduces the difficulty of the RPM problem Moreover, it is anticipated that a greater abundance of supervisory signals would correlate with enhanced intelligent performance.
However, Lico-Net+D3C still demonstrates distinct advantages that surpass other methods involving metadata. Specifically, Tranclip [28] utilizes metadata to pre-train a set of warm-start parameters for RS-Tran [28]. In other words, Tranclip fails to simultaneously endow RS-Tran with both interpretability of progressive patterns and high reasoning accuracy. Meta Triple-CFN [32] also relies on its specially designed Re-space Layer to enhance both its interpretability and reasoning accuracy. Similarly, the Re-Space Layer requires a fully trained Meta Triple-CFN to provide warm-start parameters [32]. In contrast, Lico-Net operates effectively without the necessity of a Re-space layer or a warm-start process, thereby simplifying the training procedure to a certain extent. This indicates that Lico-Net and D3C are more direct and effective, eliminating the need for indirect approaches. This ”directness” reflects a more coherent and self-consistent logic of Lico-Net and D3C compared to other methods incorporating metadata.
This paper argues that this is not an insignificant contribution, as warm-start behavior may introduce unexpected instabilities, which are not a concern for Lico-Net and D3C. For example, when Tranclip [28] uses Meta data to produce warm-start parameters for RS-Tran, early stopping is required [28]. Neglecting this step may hinder Tranclip from improving or potentially worsen the reasoning performance of RS-Tran. The Rs-space Layer [32] relies on a well-trained Meta Triple-CFN to provide warm-start parameters [32], a process that demands considerable training time and epochs. Inadequate training of these warm-start parameters can expose Meta Triple-CFN to risks akin to those faced by RS-Tran. The stopping timing for the preparation of both warm-start parameters requires manual determination and control, which brings numerous potential issues. In contrast, Lico-Net effectively mitigates these potential issues.
VIII-D Ablation experiment on D2C and D4C
In the final of this section, we devised an experiment to validate the D2C and D4C approach. Utilizing a 500-step DDPM model, we generated a substantial quantity of Bongard-Logo data, specifically 3,180 Bongard-Logo images. According to the DDPM theory, these images belong to the same distribution as the real data, albeit with unknown semantics and concepts. During the training of ResNet50 and Triple-CFN, for each feedforward pass, we introduced 14 randomly selected DDPM-generated images as auxiliary group examples for the Bongard-Logo cases. The experimental results were recorded in “D2C-” item of the Table VII.
TABLE VII: Reasoning Accuracies of models on Bongard-logo.